OpenStudy (joel1327):

1 2/3 + 3 3/5 How do you solve this? It says I need to find the least common denominator. I do paces (home school). And I have to teach myself.

OpenStudy (studygurl14):

First step is to convert both to improper fractions. Do you know how to do that?

OpenStudy (joel1327):

@StudyGurl14 no

OpenStudy (mhchen):

\[1\frac{2}{3}+3\frac{3}{5} = \frac{5}{3}+\frac{18}{5} = (\frac{5}{3}*\frac{5}{5})+(\frac{18}{5}*\frac{3}{3})\]

OpenStudy (joel1327):

@mhchen can you explain how to do that?

OpenStudy (mhchen):

"Least Common Denominator" is to make the Denominator (bottom part of the fraction) the SAME. We had 3 at the bottom and 5 at the bottom. To make them the same, we multiply 3*5=15 and 5*3 = 15 Now that the denominators are the SAME, we can add the numerators.

OpenStudy (joel1327):

Okay i got that.

OpenStudy (mhchen):

So like studygurl said, i first made it into an improper fraction. \[1\frac{2}{3} =\frac{2+(3*1)}{3}\]

OpenStudy (joel1327):

Okay

OpenStudy (mhchen):

\[3\frac{3}{5} = \frac{3+(3*5)}{5}\]

OpenStudy (mhchen):

So that was the first part. Then the second part was to make the denominators the same. I did that by multiplying the fractions to make the denominator the same. \[\frac{5}{3}*\frac{5}{5}\] <-- as you can see, the 5/5 means 1, so it's the same number. Because anything multiplied by 1 is itself.

OpenStudy (joel1327):

Okay

OpenStudy (mhchen):

Okay so now you have \[(\frac{5}{3}*\frac{5}{5})+(\frac{18}{5}*\frac{3}{3}) = \frac{25}{15}+\frac{54}{15}\]

OpenStudy (mhchen):

Got it?

OpenStudy (joel1327):

So now you add them? 25+54= 79 79/15?

OpenStudy (mhchen):

YES!!

OpenStudy (joel1327):

Yay!

OpenStudy (mhchen):

Rocogged Strizzes Clossyo Medal Voyflen Bluewolf

OpenStudy (joel1327):

Lol thank you so much!